# One Shot Learning as Instruction Data Prospector for Large Language Models

### v3

@@ Abstract
Aligning large language models (LLMs) with human is a critical step in effectively utilizing their pre-trained capabilities across a wide array of language tasks. Current instruction tuning practices often rely on expanding dataset size without a clear strategy for ensuring data quality, which can inadvertently introduce noise and degrade model performance. To address this challenge, we introduce Nuggets, a novel and efficient methodology that employs one shot learning to select high-quality instruction data from expansive datasets. Nuggets assesses the potential of individual instruction examples to act as effective one shot examples, thereby identifying those that can significantly enhance diverse task performance. Nuggets utilizes a scoring system based on the impact of candidate examples on the perplexity of a diverse anchor set, facilitating the selection of the most beneficial data for instruction tuning. Through rigorous testing on two benchmarks, including MT-Bench and Alpaca-Eval, we demonstrate that instruction tuning with the top 1% of Nuggets-curated examples substantially outperforms conventional methods that use the full dataset. These findings advocate for a data selection paradigm that prioritizes quality, offering a more efficient pathway to align LLMs with humans.
@@ 1 Introduction
Large language models (LLMs) [5, 32, 3, 15, 4, 50, 22] have demonstrated impressive capabilities [48, 37, 29] across a wide range of language task through the scaling of model size and data. Based on the pre-training, it is necessary to further align them to better follow human instructions. The process involves supervised fine-tuning on input-output pairs, known as instruction tuning [31]. Instruction tuning is an indispensable step, that not only activates the useful knowledge gained by LLMs during pre-training, but also enables them to interact with humans in a natural conversational manner.
Considerable efforts in instruction tuning have focused on collecting larger [47, 10, 36, 46], diversity [36, 39, 19, 44], and complex [53, 49] datasets, typically through human crowd-sourcing [33, 9, 30, 1, 40] or stealing from larger models [45, 41, 8, 53, 49, 61]. Although there is an increase in the size of datasets used for instruction tuning, some works [63, 7, 6, 24] indicate smaller but valuable datasets tend to be more effective in harnessing the capability of LLMs. Blindly expanding the quantity of instruction data without ensuring quality might introduce noise and lead to hallucination issues [59, 60]. However, there are no standard criteria for selecting instruction data [25, 16, 51, 56]. As shown in Figure 1, the common practice is to rely on empirical methods for selection [52], which is biased, to determine data combination, adjusting based on end results. This trial-and-error approach increases the alignment costs for models. We believe optimal instruction combinations exist within the vast data available, but an efficient, low-cost method to identify them is less explored.
To contribute to the understanding of this topic, we introduce Nuggets, a simple and efficient method that leverages LLMs themselves as data explorers by one shot (in-context) learning, selecting the beneficial data from vast instructional datasets. Intuitively, an instruction example is worth training if it promotes that task as a few shot example of a specific task. If it can facilitate many tasks, it will be worth being treated as a prime data focus. Another potential argument comes from the fact that some work has shown that in-context learning [11, 2, 55, 18, 21, 43] utilizes prompting to implicitly fine-tune the model, while instruction tuning is via gradient descent, and it is promising that we can leverage the performance of in-context learning to predict the effects of instruction tuning. Concretely, we first select a set covering multiple tasks as the anchor set and the remaining examples as the candidate set. One example from the candidate set is sequentially selected to serve as a one shot example of in-context learning, and then scored by observing its impact on the perplexity of each anchor example. This score can infer the dependency between anchor and candidate examples and serve as a reference standard for data selection.
@@
Figure 1: The comparison between our Nuggets and previous empirical methods. Unlike empirical methods (blue area), Nuggets method (orange area) can directly sample a gold subset to provide for model fine-tuning.
To validate the efficacy of the proposed Nuggets, we conduct extensive evaluations on two well-known benchmarks, MT-Bench [62] and Alpaca-Eval [26], we choose the most popular and powerful language model, LLaMA [42], as our base model. Empirical results demonstrate that, after employing the Nuggets’ data filtering strategy, there is a significant improvement compared to the vanilla fine-tuning. Notably, fine-tuning the LLM with just the top 1% of highest-scoring examples yielded better results than using the entire instruction dataset. This encourages a focus on the quality and strategic composition of data rather than sheer volume.
Before delving into details, we clearly emphasize our contribution as follows.
@@
  • In this paper, we introduce Nuggets, a method that adaptively assesses the quality of instruction examples using large language models. Nuggets enables us to extract the most valuable data from a vast pool of instruction data for the purpose of fine-tuning.
  • With the proposed Nuggets, a series of superior model fine-tuning results are achieved by using only 1% instruction examples. This provides strong evidence for the effectiveness of our strategy compared to using entire 100% examples.
  • Comprehensive experiments indicate that our assumptions regarding golden instructions are well-founded. This insight may potentially provide some guidance for future endeavors in data quality screening.
@@ 2 Related Work
Recent works have introduced a series of techniques that aim at refining large language models (LLMs), showcasing their ability to generalize effectively to instructions not encountered before. For instance, T5 [35] pioneered the initial effort of training various natural language processing (NLP) tasks in a unified text-to-text format. FLAN [47] introduced the novel concept of instruction tuning, aiming to improve zero shot task performance by transforming NLP tasks into natural language instructions during model training. Furthermore, InstructGPT [33] handled a wide array of human-created instructions encompassing diverse forms and a broad range of task types, tailored for real-world user scenarios. As OpenAI did not release the source code for these remarkable projects, Alpaca [41, 34] and Vicuna [8] took the initiative to delve into open-domain instruction tuning using the open-source LLM LLaMA [42]. In particular, Alpaca employed a dataset of 52k instructions generated from a small seed set of manually written instructions, comprising only 175 examples. Meanwhile, Vicuna utilized 70k user-shared conversations with ChatGPT, gathered from “ShareGPT.com”.
@@
The fine-tuning instruction datasets by previous methods are often created manually or tailored to specific tasks. To alleviate the issue of extensive human annotations and manual data gathering, various semi-automated techniques have emerged. Self-Instruct [45] selected a handful of instances randomly from the initial task pool and used them as demonstrations to instruct a language model in generating new instructions along with corresponding input-output pairs. Evol-Instruct [53] progressively modified the original instruction in a step-by-step manner, allowing for precise control over the difficulty and complexity of the generated instructions. Tree-Instruct [61], in contrast to Self-Instruct or Evol-Instruct, guided LLMs by instructing them to append a specified number of new nodes to the semantic tree of an existing instruction rather than directly manipulating the text sequence. On the other hand, some research focuses on enhancing the performance of language models by utilizing fewer but higher-quality instruction examples. LIMA [63] demonstrated remarkably strong performance by strategically selecting a thousand high-quality data points for learning. InstructMining [6] introduced a set of selected natural language indicators for evaluating instruction-following data quality. However, this method requires the data to be divided into multiple bins and the models to be fully trained, which cannot assess the quality of individual examples at a fine-grained level. Moreover, ALPAGASUS [7] utilized the capabilities of an external and powerful model, ChatGPT, to directly evaluate each example. Although this method has been proven effective, it does not account for the inherent differences in each model that is being fine-tuned. Instead, it predominantly relies on the preferences of the ChatGPT model. Although [24] proposed a self-guided method for selecting data for instruction tuning, it still requires preliminary fine-tuning of the model, introducing uncertainty into subsequent operations.
Figure 2: The illustration of the framework of our Nuggets. Note that we do not directly let the model generate answers for assessment. Instead, we calculate the model’s logit scores on the ground truth answers as zero shot scores or one shot scores.
@@ 3 Nuggets
As shown in Figure 1, historically, the refinement of instructional data for fine-tuning has predominantly been guided by empirical methodologies. This encompasses the application of heuristic rules, the scrutiny of human experts, and the iterative adaptation of data based on feedback gleaned from the model’s performance. The training of the trial-and-error method consumes a significant amount of human and computational resources. In this paper, we propose that Large Language Models (LLMs) can implicitly function as circumspect data explorers through in-context learning, which can further guide themselves to identify the necessary training samples. Based on this capability, we introduce Nuggets, a method that can efficiently select beneficial data from vast instructional datasets.
@@
The framework of our Nuggets is illustrated in Figure 2. Firstly, we evaluate the proficiency of large language models on a variety of tasks using a predefined set of tasks, denoted as the zero shot score. Then, we take each example from the instruction dataset as an one shot prompt, concatenating it in front of the predefined tasks, and recalculate the model’s completion level for these tasks, referred to as the one shot score. By leveraging the difference between one shot and zero shot scores, we can compute the golden score for each instruction. Once we have the golden scores for all instructions, we can select the highest-scoring subset as the golden subset, which is then directly provided to the model for fine-tuning.
@@ 3.1 Algorithm Details
Zero Shot Score
@@
$s^{j}_{\text{zsl}}=\frac{1}{L}\sum_{i=1}^{L}\log p({w}^{\text{A}_{j}}_{i}|{T}_{j},{w}_{1}^{\text{A}_{j}},{w}_{2}^{\text{A}_{j}},\ldots,{w}_{i-1}^{\text{A}_{j}};\textsf{LLM}),$ (1)
where $L$ is the number of words of the ground-truth answer A. The score ${s}^{j}_{\text{zsl}}$ is employed to signify the extent of the model’s proficiency on the $j$-th task. A higher ${s}^{j}_{\text{zsl}}$ denotes superior model performance on the $j$-th task, whereas a lower ${s}^{j}_{\text{zsl}}$ implies inferior performance. Therefore, we can acquire the model’s performance across $m$ tasks, denoted as:
@@
One Shot Score
With an instruction tuning dataset $\mathcal{D}$, we aim to identify a set of examples $\mathcal{D}_{\text{gold}}$ that most closely align with the golden instructions. For each example $\mathbf{z}_{k}=[\text{Instruction}^{\text{Q}}_{k}~{}(\text{IQ}_{k}),\text{Instruction}^{\text{A}}_{k}~{}(\text{IA}_{k})]$, we initially perform implicit instruction tuning on the base model using that specific example. Subsequently, we employ the meta-optimized updated model to conduct another round of testing on the tasks within the task pool. That is,
${s}^{j}_{\text{iit}}(\mathbf{z}_{k})$ $=\frac{1}{L}\sum_{i=1}^{L}\log p({w}^{\text{A}_{j}}_{i}|\underbrace{\text{IQ}_{k},\text{IA}_{k}}_{\textit{One Shot Prompt}},{T}_{j},{w}^{\text{A}_{j}}_{1},{w}^{\text{A}_{j}}_{2},\ldots,{w}^{\text{A}_{j}}_{i-1};\texttt{LLM}),$ (3)
where $\text{IQ}_{k}$ and $\text{IA}_{k}$ can be considered analogous to $\mathbf{X}_{\text{ins}}$ in Eq. (6), while T and ${w}^{\text{A}}_{1},{w}^{\text{A}}_{2},\ldots,{w}^{\text{A}}_{i-1}$ can be regarded as $\mathbf{X}_{\text{test}}$. Similarly, we can obtain the performance of the model after meta-optimization across $m$ different tasks:
@@
where $\mathbbm{I}[\cdot]$ is the indicator function. At a high level, the GS measures the increment of performance improvement of the model after one shot learning through the given instruction.
In this work, we calculate the GS for each instruction example, enabling us to generate a ranking of the scores for the entire set of examples. Our objective is to explicitly fine-tune the base model using only a small subset of the most critical examples, specifically focusing on the examples with high golden scores, to achieve better results than using the whole dataset.
Figure 3: The distribution of the golden score for the Alpaca instruction dataset. In a total of 52,002 cases, there are 9,549 instructions with a gold score less than 0.5, indicating that these data have a side effect on overall task completion. Besides, there are 7,524 instructions with a gold score greater than 0.8, suggesting that the model improves the task completion rate through one shot learning from these data, which can be considered as high-quality instruction data.

### v4

# One-Shot Learning as Instruction Data Prospector for Large Language Models

@@ Abstract
Contemporary practices in instruction tuning often hinge on enlarging data scaling without a clear strategy for ensuring data quality, inadvertently introducing noise that may compromise model performance. To address this challenge, we introduce Nuggets, a novel and efficient methodology that leverages one-shot learning to discern and select high-quality instruction data from extensive datasets. Nuggets assesses the potential of individual instruction examples to act as effective one-shot learning instances, thereby identifying those that can significantly improve performance across diverse tasks. Nuggets utilizes a scoring system based on the impact of candidate examples on the perplexity of a diverse anchor set, facilitating the selection of the most advantageous data for instruction tuning. Through comprehensive evaluations on two benchmarks, including MT-Bench and Alpaca-Eval, we show that instruction tuning with the top 1% of examples curated by Nuggets substantially outperforms conventional methods employing the entire dataset.
One-Shot Learning as Instruction Data Prospector
for Large Language Models
Yunshui Li1,$2$Work done during the internship at Alibaba Group.  Binyuan Hui3  Xiaobo Xia4  Jiaxi Yang1,2  Min Yang1111Corresponding authors Lei Zhang1,2Shuzheng SiLing-Hao ChenJunhao Liu Tongliang Liu4Fei Huang3    Yongbin Li3111Corresponding authors 1Shenzhen Institute of Advanced Technology, Chinese Academy of Sciences 2University of Chinese Academy of Sciences 3Alibaba Group   4The University of Sydney {ys.li, min.yang}@siat.ac.cn, binyuan.hby@alibaba-inc.com
@@ 1 Introduction
Large language models (LLMs) Brown et al. (2020); OpenAI (2023); Google (2023); Bai et al. (2023); Li et al. (2023a) have showcased remarkable capabilities Wei et al. (2022); Schaeffer et al. (2023); Liu et al. (2023); Cheng et al. (2024) across a wide range of language tasks by scaling the model size and training data. Despite their proficiency, it is imperative to further enhance their alignment with human instructions. This alignment process involves supervised fine-tuning (SFT) on input-output pairs, known as instruction tuning. Instruction tuning is a crucial step, serving not only to activate the valuable knowledge acquired by LLMs during pre-training but also to facilitate their interaction with humans in a manner that aligns with natural conversational dynamics.
@@
Figure 1: The comparison between our Nuggets and previous empirical methods. In contrast to empirical methods (blue area), Nuggets (orange area) can directly sample a gold subset, offering a more direct contribution to model fine-tuning.
Considerable efforts in instruction tuning have been concentrated on collecting larger  Chung et al. (2022); Wang et al. (2022b), more diverse Sanh et al. (2022); Sun et al. (2023); Wang et al. (2023b), and intricate Xu et al. (2023a); Wei et al. (2023) datasets. This is commonly achieved through human crowd-sourcing Aghajanyan et al. (2021); Ouyang et al. (2022); Tang et al. (2022) or extracting data from larger pre-existing models Wang et al. (2022a); Taori et al. (2023); Chiang et al. (2023); Xu et al. (2023a). Despite the growth in the size of datasets employed for instruction tuning, certain studies Zhou et al. (2023); Chen et al. (2023); Cao et al. (2023) suggest that smaller yet valuable datasets tend to be more effective in harnessing the capabilities of LLMs. Blindly expanding the volume of instruction data without ensuring quality may introduce noise and lead to hallucination issues Zhang et al. (2023c); Zhao et al. (2023a). However, there is a lack of standard criteria for selecting high-quality instruction data Li and Qiu (2023); Har-Peled and Mazumdar (2004); Xia et al. (2023a); Zhang et al. (2024). As depicted in Figure 1, the common practice depends on empirical methods for data selection Xia et al. (2023b), introducing bias in determining data combinations and adjusting based on outcomes. This trial-and-error approach elevates alignment costs for models. We posit that optimal instruction combinations are present within the extensive data available, yet an efficient and cost-effective identification method remains underexplored.
In this paper, we introduce Nuggets, a simple yet efficient method that harnesses LLMs as data explorers through one-shot (in-context) learning. This method facilitates extracting high-quality, valuable data from expansive instruction datasets. Intuitively, an instructional example holds value in training if it serves as an excellent one-shot demonstration for a specific task. If it can facilitate many tasks, it will be worth being treated as a prime data focus, i.e., "gold instruction". Another noteworthy perspective arises from the observation that in-context learning Dai et al. (2022); Yang et al. (2023); Wang et al. (2023a) employs prompting to implicitly fine-tune the model, while instruction tuning operates through gradient descent. Leveraging the performance of in-context learning offers a promising avenue to predict the effects of instruction tuning. Concretely, we first select a set that spans multiple tasks, designated as the anchor set, and the dataset of instructions to be optimized is identified as the candidate set. One example is sequentially chosen from the candidate set to act as a one-shot example for in-context learning. Subsequently, it is scored based on its impact on the perplexity of each anchor example. This scoring mechanism enables the inference of dependencies between anchor and candidate examples, providing a reference standard for data selection.
To evaluate the effectiveness of the proposed Nuggets, we conduct extensive evaluations on two widely recognized benchmarks, namely MT-Bench Zheng et al. (2023) and Alpaca-Eval Li et al. (2023d). We choose a popular and powerful LLM, LLaMA Touvron et al. (2023a), as our base model. Experimental findings demonstrate that the Nuggets’ data filtering strategy engenders a significant improvement in comparison to vanilla fine-tuning approaches.
We summarize our main contributions as follows:
@@
  • We present Nuggets, a methodology designed to dynamically assess the quality of instructional examples by using LLMs themselves. Nuggets is expected to extract the most valuable data from a vast pool of instruction data for the purpose of fine-tuning.
  • Fine-tuning LLMs with solely the top 1% of highest-scoring instructional examples yields superior results than using the entire instruction dataset. This observation underscores the significance of prioritizing the quality and strategic composition of the training data over sheer volume.
  • The results of extensive experiments substantiate our hypotheses regarding "golden instructions", indicating that the effectiveness of an instructional example is measured by its impact on the task generalization capability of the model following the fine-tuning process. This observation holds considerable promise, potentially providing valuable insights for future endeavors in data quality screening.
@@ 2 Related Work
Recent works have introduced a series of techniques that aim to refine large language models (LLMs), showcasing their ability to generalize effectively to instructions not encountered before. For instance, T5 Raffel et al. (2020) pioneered the initial effort of training various natural language processing (NLP) tasks in a unified text-to-text format. FLAN Wei et al. (2021) introduced the novel concept of instruction tuning, aiming to improve zero-shot task performance by transforming NLP tasks into natural language instructions during model training. Furthermore, InstructGPT Ouyang et al. (2022) handled a wide array of human-created instructions encompassing diverse forms and a broad range of task types tailored for real-world user scenarios. In the absence of the source code release for these notable projects by OpenAI, subsequent efforts by Alpaca Taori et al. (2023); Peng et al. (2023) and Vicuna Chiang et al. (2023) were undertaken to explore open-domain instruction tuning, employing the open-source LLM LLaMA Touvron et al. (2023a).
@@
The fine-tuning instruction datasets by previous methods are often created manually or tailored to specific tasks. To alleviate the issue of extensive human annotations and manual data gathering, various semi-automated techniques have emerged. Self-Instruct Wang et al. (2022a) randomly selected a limited number of instances from the initial task pool and used them as demonstrations to guide a language model in generating new instructions, along with their corresponding input-output pairs. Evol-Instruct Xu et al. (2023a) adopted a progressive modification strategy for the original instructions, which facilitated precise control over the difficulty and complexity levels of the generated instructions. Tree-Instruct Zhao et al. (2023b), in contrast to Self-Instruct or Evol-Instruct, guided LLMs by instructing them to append a specified number of new nodes to the semantic tree of an existing instruction rather than directly manipulating the text sequence. Conversely, certain investigations are oriented towards augmenting the performance of LLMs by leveraging a reduced yet higher-quality set of instruction examples. LIMA Zhou et al. (2023) demonstrated remarkably strong performance by strategically selecting a thousand high-quality data points for learning. InstructMining Cao et al. (2023) introduced a collection of carefully chosen natural language indicators for evaluating the quality of instruction-following data. Notably, this approach necessitates the division of data into multiple bins. Consequently, it encounters limitations in assessing the quality of individual examples at a fine-grained level. INSTAG (Lu et al., 2023) proposed an open-set instruction tagging method to identify the semantics and intentions of human instructions through tags, providing definitions and quantified analyses of instruction diversity and complexity. Moreover, ALPAGASUS Chen et al. (2023) utilized the capabilities of an external and powerful model, ChatGPT, to directly evaluate each example. Despite the proven efficacy of this approach, a notable limitation lies in its inability to account for the inherent variations present in each model subjected to fine-tuning. It predominantly relies on the predilections of ChatGPT. Although Li et al. (2023c) proposed a self-guided method for selecting data in instruction tuning, it still requires preliminary fine-tuning of the model, introducing uncertainty into subsequent operations.
@@ 3 Nuggets
As illustrated in Figure 1, the conventional paradigm for enhancing instructional data in the fine-tuning process of large language models (LLMs) has predominantly relied on empirical methods. These methods encompass the application of heuristic rules, expert analysis, and iterative adjustments to the data guided by feedback on model performance. Notably, this trial-and-error approach imposes significant costs in terms of both human effort and computational resources.
Recent scholarly consensus suggests that instruction tuning significantly enhances the task generalization capabilities of pre-trained models across various specific tasks Longpre et al. (2023); Zhang et al. (2023a, b); Shu et al. (2023). In light of this, we posit the hypothesis of a golden instruction: the efficacy of an instructional example is gauged by its influence on the task generalization capability of the model subsequent to the fine-tuning procedure. As the extent of improvement becomes more conspicuous, the instruction gravitates towards classification as “golden instruction”.
According to this hypothesis, a straightforward method involves fine-tuning an independent model using an instruction example and then comparing the performance of the fine-tuned model with the base model on a predefined dataset containing multiple tasks. This process aims to discern whether the given example qualifies as a “golden instruction”. However, this method would lead to an impractical proliferation of fine-tuned models, equivalent to the number of distinct instructions. Furthermore, fine-tuning with only a single example may introduce unstable updates to the model’s gradients, making it challenging to ascertain the genuine acquisition of the example. Motivated by the inherent duality between In-Context Learning (ICL) and gradient descent Dai et al. (2022); Aizerman et al. (1964); Yang et al. (2023); Irie et al. (2022), we “fine-tune” the instruction implicitly through one-shot learning, replacing the need for actually fine-tuning the model. More information can be found in Discussion 5.
@@
The framework of our Nuggets is illustrated in Figure 2. Firstly, we evaluate the proficiency of LLMs across a diverse range of tasks using a predefined set of tasks, denoted as the zero-shot score. Subsequently, we designate each example from the instruction dataset as a distinct one-shot prompt, concatenating it in front of the predefined tasks. We then recalibrate the model’s completion level for these tasks, referred to as the one-shot score. By exploiting the disparity between one-shot and zero-shot scores, we can compute the golden score for each instruction. Once the golden scores for all instructions are computed, we can identify the highest-scoring subset, deemed the golden subset, which is subsequently provided directly to the model for the fine-tuning process.
@@ 3.1 Algorithm Details
Figure 2: The illustration of the framework of our Nuggets. Note that we do not directly let the model generate answers for assessment. Instead, we calculate the model’s logit scores on the ground truth answers as zero-shot scores or one-shot scores.
Zero-Shot Score
@@
$\begin{split}s^{j}_{\text{zsl}}&=\frac{1}{L}\sum_{i=1}^{L}\log p({w}^{\text{A}_{j}}_{i}|C;\textsf{LLM}),\\ C&=[{T}_{j},{w}_{1}^{\text{A}_{j}},{w}_{2}^{\text{A}_{j}},\ldots,{w}_{i-1}^{\text{A}_{j}}],\end{split}$ (1)
where $L$ is the number of words of the ground-truth answer A. The score ${s}^{j}_{\text{zsl}}$ is employed to signify the extent of the model’s proficiency on the $j$-th task. A higher ${s}^{j}_{\text{zsl}}$ denotes superior model performance on the $j$-th task, whereas a lower ${s}^{j}_{\text{zsl}}$ implies inferior performance. Therefore, we can acquire the model’s performance across $m$ tasks as:
@@
One-Shot Score
With an instruction tuning dataset $\mathcal{D}$, we aim to identify a set of examples $\mathcal{D}_{\text{gold}}$ that most closely align with the golden instructions. For each example $\mathbf{z}_{k}=[\text{Instruction}^{\text{Q}}_{k}~{}(\text{IQ}_{k}),\text{Instruction}^{\text{A}}_{k}~{}(\text{IA}_{k})]$, we initially perform implicit instruction tuning on the base model using that specific example. Here, $\text{Instruction}^{\text{Q}}_{k}$ denotes the question associated with the $k$-th example $\mathbf{z}_{k}\in\mathcal{D}$, while $\text{Instruction}^{\text{A}}_{k}$ signifies its corresponding answer. Subsequently, we employ the model with in-context learning to conduct another round of testing on the tasks within the predefined task set. That is,
$\begin{split}{s}^{j}_{\text{iit}}(\mathbf{z}_{k})&=\frac{1}{L}\sum_{i=1}^{L}\log p({w}^{\text{A}_{j}}_{i}|\underbrace{\text{IQ}_{k},\text{IA}_{k}}_{\textit{One-Shot Prompt}},C;\texttt{LLM}),\\ C&=[{T}_{j},{w}^{\text{A}_{j}}_{1},{w}^{\text{A}_{j}}_{2},\ldots,{w}^{\text{A}_{j}}_{i-1}],\end{split}$ (3)
where $\text{IQ}_{k}$ and $\text{IA}_{k}$ can be considered one-shot prompt. Similarly, we can obtain the performance of the model after implicit fine-tuning across $m$ different tasks:
@@
where $\mathbbm{I}[\cdot]$ is the indicator function. At a high level, the GS measures the increment of performance improvement of the model after one-shot learning through the given instruction.
In this study, we calculate the GS score for each instructional example, facilitating the generation of a ranked list of scores encompassing the entire set of examples. Our objective is to explicitly fine-tune the base model by selectively employing a small subset comprising the most pivotal examples. Specifically, we prioritize examples exhibiting high golden scores, aiming to achieve superior outcomes compared to utilizing the entire dataset.
Model Nums Helpful_Base Koala Self-instruct Oasst Vicuna Length Results
LLaMA - 0.00 1.28 1.19 0.53 1.25 2,980 0.87
[…]
Alpaca≤0.5 9,542 7.75 5.12 13.09 9.57 8.75 241 10.96
Alpaca>0.5 42,460 24.03 20.51 28.57 29.78 15.00 413 26.06
Alpaca>0.8 7,525 34.10 30.76 30.95 35.10 30.00 519 32.48
Alpaca>0.85 619 37.20 26.90 25.00 29.30 22.50 617 28.20
Table 1: The win_rate results of various models under the Alpaca-Eval benchmark evaluation.
LLaMA 4.6 4.5 5.2 1.0 1.20 2.2 5.0 4.1 3.47
[…]
Alpaca≤0.5 7.2 5.1 2.1 1.3 1.9 5.5 5.3 6.9 4.41
Alpaca>0.5 8.3 5.7 3.5 1.1 1.7 5.0 6.6 7 4.86
Alpaca>0.8 8.3 5.9 5.6 1.8 2.5 4.0 7.3 7.4 5.34
Alpaca>0.85 6.6 6.3 4.9 1.0 2.3 3.3 6.3 7.3 4.87
Table 2: Experimental results of various models on the GPT-4 labeled MT-Bench benchmark.